OpenStudy (anonymous):

Solve for x. Show all work please. 3(x + 2) = -5 - 2(x - 3)

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

Then you should be able to do that on your own.

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

uhhuh, sure

OpenStudy (anonymous):

Anyway I'll do it for you.

OpenStudy (anonymous):

3x+6 = -5 -2x+6

OpenStudy (anonymous):

could someone explain?

OpenStudy (anonymous):

subtract 6

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

3x=-5-2x

OpenStudy (texaschic101):

take the number right before the parenthesis and multiply it by everything in the parenthesis 3(x + 2) = -5 -2 (x - 3) 3*x + 3*2 = -5 - 2*x - 2 * -3 now just work this out and what do you get

OpenStudy (texaschic101):

you gonna try ?

OpenStudy (anonymous):

i am but its still fuzzy

OpenStudy (texaschic101):

3*x + 3*2 = -5 - 2*x - 2*-3 3x + 6 = -5 - 2x + 6 do you understand how I got that ?

OpenStudy (anonymous):

i think i do but is this how i would get the answer like is -5 - 2*x -2*-3

OpenStudy (anonymous):

the answer? sorry for the lag

OpenStudy (anonymous):

3(x + 2) = -5 - 2(x - 3) 3x+6=-5-2x+6 3x+6=1-2x 3x+2x=1-6 5x=-5 x=-5/5 x=-1

OpenStudy (texaschic101):

I am just showing steps in how you multiply the number right before the parenthesis by everything in the parenthesis. By doing this, the parenthesis are removed. for instance : 3 (x + 2) = 3*x + 3*2 = 3x + 6 -2(x - 3) = -2*x - 2*-3 = -2x + 6 understand ?

OpenStudy (anonymous):

is it helpful for u manguitar..??

OpenStudy (texaschic101):

vanshunegi just posted the answer and it is correct